2. Two 5 ohm resistors are connected in series to a 15 V battery. What current flows through the circuit? a. 0.33 Amperes b. 66 Amperes C. 1.500 Amperes d. 00 Amperes e. 150 Amperes
Answer #1

total resistance= R=5+5= 10 ohm

V= 15 volt

V=I R

i= V/R

i= 15/10

i= 1.5 A.

so the current = 1.5 amperes